OpenStudy (anonymous):

prove that: sinxtanx/1-cosx is congruent to 1 + 1/cosx

OpenStudy (anonymous):

\[\frac{ \sin x~\tan x }{ 1-\cos x }\times \frac{ 1+\cos x }{ 1+\cos x }=?\]